# Comparative Study of Recurrent Neural Networks for Virtual Analog Audio Effects Modeling

### v3

@@ 3.3 Audio Effects and Datasets
In this section, we introduce the audio effects selected for the study and the composition of their datasets. We proceeded to record the dataset using effects that were at our disposal. In particular, we selected the following effects: overdrive, saturation, equalization, low-pass filter, and compression. Table 1 summarizes all the datasets included in the study. Delay-based effects require a different approach than a direct black box  [49],  [48]. For this reason, they are not included in this study.
The dataset we use is derived from the following hardware devices: Behringer OD300 overdrive pedal  [^3] https://www.behringer.com/product.html?modelCode=P0608, Behringer Neutron  [^4] https://www.behringer.com/product.html?modelCode=P0CM5’s overdrive module, Behringer Neutron’s filter module in low-pass mode, TC Electronic Bucket Brigade Analog Delay used as a saturator  [^5] https://www.tcelectronic.com/product.html?modelCode=P0EBV, CL 1B TubeTech  [^6] http://www.tube-tech.com/cl-1b-opto-compressor/and Teletronix LA-2A  [^7] https://www.uaudio.com/hardware/la-2a.html optical compressors. We have also collected data from two software VA plugins: the Helper Saturator  [^8] https://www.waproduction.com/plugins/view/helper-saturator and Universal Audio Pultec EQ  [^9] https://www.uaudio.com/uad-plugins/equalizers/pultec-passive-eq-collection.html, because the hardware saturator presents no variable control parameter, and because we have no analog equalizer in our immediate availability.
@@
Both the overdrive devices feature the following control parameters: volume, distortion level, and tone. The OD300 is a distortion pedal typically used with electric guitars, which features two different types of distortion, labeled overdrive and distortion modes. To record the dataset the unit was set to overdrive mode. The device was sampled at $5$ different and equally spaced values for the distortion and tone knobs, between the minimum and maximum values, resulting in a $25$ different combination. A separate output signal signal is recorded for each combination of the control parameters. Neutron’s Overdrive is a module integrated into the Neutron synthesizer. Thanks to the semi-modular nature of this device, we were able to isolate the overdrive module. We recorded the dataset’s varying distortion and tone as we did for the OD300, but sampling $10$ equally spaced values for each parameter. In this case, we recorded the output for $100$ different combinations of control parameters. In both devices, the volume level controls an attenuator, and we fixed this parameter to the maximum (i.e. no attenuation).
@@
Similarly to the overdrive, the Behringer Neutron integrates a filter module, which features frequency cutoff and resonance knobs. The filter can be set as low-pass, high-pass, and band-pass. We isolated the module and recorded the dataset at $10$ equally spaced values for the cutoff and $5$ for the resonance. The cutoff parameter is between the minimum and maximum, while the resonance parameter is between the minimum to the mid value. having $100$ different combinations and selecting the low-pass filter type. The filter was disconnected from the ADSR envelope module.
@@
Helper Saturator software emulates analog saturation, in particular tube and tape saturation. The software presents four parameters: lowpass and highpass cutoffs, saturation type switch (tape or tube), and saturation level. We recorded the dataset for $10$ different values of the saturation level between minimum and maximum. The saturation switch was set to tape mode. Lowpass and highpass cutoffs were set to $20$ and $20000$ Hz, respectively.
The other device under this category is the TC Electronic Bucket Brigade Analog Delay pedal. The pedal is equipped with the well-known bucket brigade circuit, which consists of a series of capacitors that transmit the signal. By regulating the speed at which the capacitors fill and empty, a clock controls and slows down the signal, resulting in a delay as it passes through the bucket brigade device (BBD) circuit. The signal is sampled by storing an electric charge in a capacitor at each clock pulse. The sampling creates harmonic content and saturation. After traversing the BBD, the signal undergoes another low-pass filtering process, attenuating the new harmonics, and is expanded to match the dynamic range of the original input signal as closely as possible. Before entering the BBD section, the signal is typically manipulated as well, compressing and filtering using an anti-aliased filter.
This device features four knobs: time, feedback, volume, and depth, as well as a mode switch. The time knob determines the length of the delay, adjustable up to $600$ ms. The feedback knob controls the number of repetitions of the delayed signal. The volume knob adjusts the volume of the delayed signal, while the depth knob defines the level of modulation. Activating the mode switch enables the modulation effect.
For our study, we primarily want to capture the processed signal without the delay component, which can be digitally simulated  [60]. All knobs were set to their minimum values except for the volume knob, while the mode switch remained deactivated.
@@
The software plug-in Universal Audio Pultec Passive EQ emulates the tube-based equalizer Pultec EQP-1A. The effect presents low-frequency ([20, 30, 60, 100] Hz) and high-frequency ([3, 4, 5, 8, 10, 12, 16]kHz) switches to select the frequencies to boost or cut using the respective boost and attenuator knobs ([0,10]). Additionally, the emulation also has a bandwidth knob ([0,10]). The equalizer boosts and cuts the selected frequency bands, giving some saturation simultaneously. In this case, we recorded the dataset varying the boost and attenuator knobs at $5$ different equally-spaced values up to half of their range, setting $60$ Hz as the low-frequency switch value and 00 (sharp) as the bandwidth, obtaining $25$ different recordings.
@@
The CL 1B and the LA-2A are analog optical compressors. In optical compressors, a lighting-emitting element is fed with the audio signal that illuminates a light-sensitive resistor. The input signal’s amplitude determines the element’s brightness, which, in turn, changes the resistance in the gain attenuation circuit. We recorded the dataset directly from the compressor. The CL 1B compressor dataset presents four parameters: attack, release, ratio, and threshold. We consider the threshold values to range from 00 to $40$ dBu, ratio values from $2$:$1$ to $10$:$1$, attack time from $0.5$ to $300$ ms, and release time from $0.05$ to $10$ seconds. We consider three equally spaced values for the ratio, attack, and release time and four values for the threshold. The resulting number of parameter combinations is $108$. On the other hand, LA-2A presents the peak reduction knob indicating the amount of compression ([00,$100$]) and a mode switch (limiter, compression). We consider ten equally spaced values for the peak reduction, resulting in $20$ parameter combinations.
@@
OD300111Hardware. Overdrive Level [min, max] $25$
@@
Neutron’s Overdrive Level [min, max] $100$
OD module [^1] Hardware. Tone [min, max]
Neutron’s Low pass Cutoff [min, max] $100$
Filter module [^1] Hardware. filter Resonance [min, max]
@@
Saturator [^2] Software. Saturation Type Tape
@@
Passive EQ  [^2] Software. Frequency booster [0,10]
@@
LA-2A [^1] Hardware.  [^3] Downloaded. Optical Peak Reduction [0,100] $20$
@@
CL 1B [^1] Hardware. Optical Threshold [-40, 0] dB $108$
@@ 3.3.1 Data Collection
The data collection process was carried out using a MOTU M4 audio interface to feed a selection of audio signals into the system and simultaneously record its output. To achieve this, the left input channel of the audio interface was connected to the left output channel of the interface itself. The right input channel of the audio interface was connected to the output of the device, and the input of the device to the right output channel of the audio interface. This allows recording effectively both the device’s input and output signals compensating for the minor sound coloring and latency of the audio interface. The audio data was recorded at a sampling rate of $48$ kHz.
@@ 3.4 Experimenting and Learning
The models are trained for $60$ epochs and using the Adam [61] optimizer with a gradient norm scaling of $1$  [62]. The training was stopped earlier in case of no reduction of validation loss for $10$ epochs. We design a time-based schedule for the learning rate as follows:
$lr=\frac{LR}{e^{4}},\text{ with }LR=3\cdot 10^{4}$ (13)
where LR is the initial learning rate, and $e$ is the number of the epochs. Test losses and evaluation metrics are computed using the model’s weights that minimize the validation loss throughout the training epochs. Finally, the input signal is split into segments of $2400$ samples (equivalent to $50$ ms), to be processed before updating the weights.
We adopted the same learning schedule and minibatch size for all models across all datasets. As explained in Section 3.3, we collected the datasets using the same approach and the same input file to minimize the difference among them. However, it is important to highlight that each dataset and modeling task may benefit from specific adjustments to achieve more accurate results. Different datasets and tasks may require slightly different learning rate schedules; an excessively slow or fast learning rate can lead to suboptimal solutions. The number of samples to process before updating the weights, referred to as the minibatch size, impacts the frequency of updating weights: larger sizes may result in faster convergence but poorer generalization, while smaller sizes can aid generalization but may require additional epochs.
@@ 4 Results
The evaluation is based on the metrics described in Section 3.2 and presented in Table 3, and 4. The MSE loss and the epoch-related epoch resulting in lower validation loss during training are reported in Table 2. The losses and the metrics consider all conditioning scenarios and provide an average error across all combinations of parameters. To further verify the model’s capacity to learn the conditioned effect, we provide a series of plots that compare the target signal with the predictions generated by the four distinct architectures. For these plots, we chose the modeling scenario that we believe to be the most challenging based on the effect parameters used as conditioning factors for the networks. Except for the low-pass filter, this typically involves parameters at the maximum of the selected range since the output signal, under these conditions, is likely to deviate most significantly from the input. Regarding the low-pass filter, the parameters are configured for minimum cutoff and maximum resonance. The plots show the RMSE in Figure  4, the spectral flux in Figure  5, the spectra in Figure  7, and the spectrograms in Figure  6. For datasets including conditioning parameters, the plots showing predictions versus the target are generated for a $2.5$ seconds input signal containing bass, guitar, and drum loop sound. These input signals are part of the respective test sets. To ensure a clearer comparative visualization of the spectral flux, we use only the first $1.25$ seconds.
In the OD300 overdrive dataset, the LSTM model presents the lowest MSE, MAE, and ESR, indicating better general amplitude matching between target and prediction. The S4D model is also accurate, especially when considering the transients and the RMS envelope. This can be observed from the lower error in terms of RMSE and spectral flux. On the other hand, in the case of maximum distortion, as seen in Figures 4 and 5, this superiority is less perceived. ED and LRU are left behind except when looking $M_{SF}$ where ED show better performance than LSTM. Notably, the spectrograms shown in Figure 6 illustrate that S4D model generates more harmonics in the higher frequency region than LSTM, and especially than LRU whose produces a cleaner signal. This could be perceptually advantageous compared to the other model that poorly predicts such region of the spectrum, as summarized by the $M_{MFCC}$ errors. From Figure 7 it is evident how the LSTM-based produces the best matching for energy at low frequencies.
Figure 5: Comparison of the spectral flux between the target and predicted output. The results refer to the highest conditioning values, representing the most challenging scenario.
Models trained on the Neutron OD dataset exhibit larger errors, suggesting that this dataset presents a more challenging modeling case for the architectures being investigated. Both LSTM and LRU models yield similar results in terms of average amplitude and frequency errors. The LSTM models show slightly better performance in the time domain, while the LRU excel in the frequency domain. As observed in Figure 4, both LSTM and LRU model achieves a better match with the target. However, the LSTM model tends to produce more energy, whereas the LRU model yields less energy than the target. The S4D model surpasses previous models in terms ofSF, despite exhibiting a larger mismatch at maximum distortion settings, as visible in Figure  5. The spectrograms indicate that the S4D model exhibits a slightly better match than the LSTM model, even though the latter exhibits smaller errors. Upon informal listening, the S4D model tends to add less distortion necessary, whereas the LSTM-based models tend to do the opposite. Lastly, the LRU model exhibits a later stopping epoch than the other three cases.
For models trained on the Saturator dataset, the LRU model presents lower MSE, MAE, and $M_{RMSE}$ values. Meanwhile, the S4D model outperforms the others with respect to ESR and $M_{SF}$ metrics. The ED model is slightly less accurate but performs similarly to S4D. In terms of frequency domain errors, LRU also produces better results. As evidenced in Figures  4 and  5, the S4D and ED models achieve closer matching to the target, whereas the LRU model exhibits significantly lower energy, and the LSTM model introduces artifacts, which can be clearly heard when listening to the audio output and negatively affect its performance. These artifacts can also be seen in Figure 5 and  6. Despite these issues, the S4D and ED models still provide a better match with the target spectrograms. The same trend is observed in Figure 7, suggesting that the LRU model performs better in scenarios where minimal sound alteration is required.
[…]
Figure 7: Comparison of the frequency content between the target and predicted output. The FFT is computed using windows of $512$ samples. The results refer to the highest conditioning values, representing the most challenging scenario.
The optical compressor features a time-variant response with pronounced temporal dependencies, largely influenced by its variable attack and release parameters. This is especially noticeable in the case of the CL 1B compressor, which represents a more challenging task for neural network modeling. Among the tested models, the LSTM is the least accurate, exhibiting higher loss in both the LA-2A and CL 1B cases. In contrast, the S4D and ED models demonstrate superior accuracy for these tasks. Specifically, S4D outperforms the ED in the LA-2A case, while the textbfED excels in the CL 1B case when evaluated using time domain-based metrics. However, the trend reverses in the frequency domain, although ED consistently shows lower $M_{SF}$ and $M_{MFCC}$ values for both cases.
In the LA-2A case, we observe an overall good match for RMSE in Figure 4 and for transient phases in Figure 5. The LSTM and LRU models display slightly more mismatches in both plots. Upon examining Figure  6, spectrograms related to the LRU model appear similar to the target output than those of the S4D model, despite a discernible decrease in accuracy observed in Figure 7. In the case of the compressor, the capability to encode past information of the signal is critical, and S4D exhibits overall better accuracy. The ED and LRU models are designed to improve the temporal tracking abilities of basic LSTM, as confirmed by the results for both compressor devices. This improvement is particularly evident with the CL 1B when setting long attack and release time through the conditioning control parameters. These reach up to $300$ ms and $5$ seconds, respectively. In Figure 4 and  5, the difference across models is subtler, yet the LRU model consistently shows smaller RMS errors than other models and a more accurate match in the frequency domain as visible in Figure  6 and  7.
Modeling the low-pass filter is inherently a more complex task than the equalizer, given its more pronounced alterations to the frequency content of the processed signal. Nevertheless, informal listening evaluations of the predicted outputs indicate that all models struggle with accurately emulating the filter behavior; instead of eliminating high-frequency contents, they attenuate the whole signal, as apparent in Figure 6. However, the LRU model outperforms the others when considering the quantitative metrics. The LSTM model, in particular, tends to incorrectly predict excessive energy in the lower spectrum, as shown in Figure 7.
@@
OD300 LSTM $\boldsymbol{2.46\cdot 10^{-4}}$ $144$ Neutron’s LSTM $\boldsymbol{2.27\cdot 10^{-3}}$ $4$
ED $3.57\cdot 10^{-4}$ $176$ OD ED $2.79\cdot 10^{-3}$ $3$
@@
LA-2A LSTM $1.11\cdot 10^{-4}$ $59$ CL 1B LSTM $1.78\cdot 10^{-4}$ $74$
ED $9.54\cdot 10^{-5}$ $90$ ED $\boldsymbol{8.96\cdot 10^{-5}}$ $38$
LRU $1.14\cdot 10^{-4}$ $146$ LRU $6.76\cdot 10^{-4}$ $168$
S4D $\boldsymbol{8.43\cdot 10^{-5}}$ $62$ S4D $1.22\cdot 10^{-4}$ $14$
@@
LRU $\boldsymbol{4.14\cdot 10^{-4}}$ $176$ EQ LRU $1.03\cdot 10^{-3}$ $150$
@@
Neutron’s LSTM $4.07\cdot 10^{-3}$ $46$

### v5

@@ 3.3 Audio Effects and Datasets
In this section, we introduce the audio effects selected for the study and the composition of their datasets. We proceeded to record the dataset using effects that were at our disposal. In particular, we selected the following effects: overdrive, saturation, equalization, low-pass filter, and compression. Table 2 summarizes all the datasets included in the study. Delay-based effects require a different approach than a direct black box  (Wright & Välimäki,, 2021),  (Mikkonen et al.,, 2023). For this reason, they are not included in this study.
The dataset we use is derived from the following hardware devices: Behringer OD300 overdrive pedal  [^2] https://www.behringer.com/product.html?modelCode=P0608, Behringer Neutron  [^3] https://www.behringer.com/product.html?modelCode=P0CM5’s overdrive module, Behringer Neutron’s filter module in low-pass mode, TC Electronic Bucket Brigade Analog Delay used as a saturator  [^4] https://www.tcelectronic.com/product.html?modelCode=P0EBV, CL 1B TubeTech  [^5] http://www.tube-tech.com/cl-1b-opto-compressor/and Teletronix LA-2A  [^6] https://www.uaudio.com/hardware/la-2a.html optical compressors. We have also collected data from two software VA plugins using  (Fasciani et al.,, 2024b): the Helper Saturator  [^7] https://www.waproduction.com/plugins/view/helper-saturator and Universal Audio Pultec EQ  [^8] https://www.uaudio.com/uad-plugins/equalizers/pultec-passive-eq-collection.html, because the hardware saturator presents no variable control parameter, and because we have no analog equalizer in our immediate availability.
@@
Both the overdrive devices feature the following control parameters: volume, distortion level, and tone. The OD300 is a distortion pedal that features two different types of distortion, labeled overdrive and distortion modes. To record the dataset, the unit was set to overdrive mode. The device was sampled at $5$ different and equally spaced values for the distortion and tone knobs. Neutron’s overdrive is a module integrated into the Neutron synthesizer. Thanks to the semi-modular nature of this device, we were able to isolate the overdrive module. We recorded the dataset’s varying distortion and tone as we did for the OD300, but sampling $10$ equally spaced values for each parameter. In both devices, the volume level controls an attenuator, and we fixed this parameter to the maximum (i.e., no attenuation).
@@
Similarly to the overdrive, the Behringer Neutron integrates a filter module, which features frequency cutoff and resonance knobs. The filter can be set as low-pass, high-pass, and band-pass. We isolated the module and recorded the dataset at $10$ equally spaced values for the cutoff and $5$ for the resonance, selecting the low-pass filter type. The filter was disconnected from the ADSR envelope module.
@@
Helper Saturator software emulates analog saturation, in particular tube and tape saturation. The software presents four parameters: lowpass and highpass cutoffs, saturation type switch (tape or tube), and saturation level. We recorded the dataset for $10$ different values of the saturation level. The saturation switch was set to tape mode. Lowpass and highpass cutoffs were set to $20$ and $20000$ Hz, respectively.
@@
The software plug-in Universal Audio Pultec Passive Equalizer emulates the tube-based equalizer Pultec EQP-1A. The effect presents low-frequency ([20, 30, 60, 100] Hz) and high-frequency ([3, 4, 5, 8, 10, 12, 16]kHz) switches to select the frequencies to boost or cut using the respective boost and attenuator knobs ([0,10]). Additionally, the emulation also has a bandwidth knob ([0,10]). The equalizer boosts and cuts the selected frequency bands, giving some saturation simultaneously. In this case, we recorded the dataset varying the boost and attenuator knobs at $5$ different equally-spaced values up to half of their range, setting $60$ Hz as the low-frequency switch value and 00 (sharp) as the bandwidth.
@@
The CL 1B and the LA-2A are analog optical compressors. In optical compressors, a lighting-emitting element is fed with the audio signal that illuminates a light-sensitive resistor. The input signal’s amplitude determines the element’s brightness, which, in turn, changes the resistance in the gain attenuation circuit. We recorded the dataset directly from the compressor. The CL 1B compressor dataset presents four parameters: attack, release, ratio, and threshold. We consider three equally spaced values for the ratio, attack, and release time and four values for the threshold. On the other hand, LA-2A presents the peak reduction knob indicating the amount of compression ([00,$100$]) and a mode switch (limiter, compression). We consider 10 equally spaced values for the peak reduction.
@@
OD300111Recorded from hardware devices. Overdrive Level [min, max] $25$
@@
Neutron Overdrive Level [min, max] $100$
Overdrive module [^1] Recorded from hardware devices. Tone [min, max]
Neutron Low pass Cutoff [min, max] $100$
Filter module [^1] Recorded from hardware devices. filter Resonance [min, max]
@@
Saturator [^2] Recorded from software devices. Saturation Type Tape
@@
Passive EQ  [^2] Recorded from software devices. Frequency booster [0,10]
@@
LA-2A [^1] Recorded from hardware devices.  [^3] Publicly available. Optical Peak Reduction [0,100] $20$
@@
CL 1B [^1] Recorded from hardware devices. Optical Threshold [-40, 0] dB $108$
@@ 3.3.1 Data Collection
The data collection process was carried out using a MOTU M4 audio interface to feed a selection of audio signals into the system and simultaneously record its output. To achieve this, the left input channel of the audio interface was connected to the left output channel of the interface itself. The right input channel of the audio interface was connected to the output of the device, and the input of the device to the right output channel of the audio interface. This allows effective recording of both the device’s input and output signals, compensating for the minor sound coloring and latency of the audio interface. The audio data was recorded at a sampling rate of $48$ kHz.
@@ 3.4 Experimenting and Learning
The models are trained for $200$ epochs and using the Adam (Kingma & Ba,, 2014) optimizer with a gradient norm scaling of $1$  (Pascanu et al.,, 2013). The training was stopped earlier in case of no reduction of validation loss for $10$ epochs. We design a learning rate with an exponential decay according to the formula: $lr=LR*0.25^{e}$, where LR is the initial learning rate and it is set to $3\cdot 10^{-4}$, and $e$ is the number of the epochs. Test losses and evaluation metrics are computed using the model’s weights that minimize the validation loss throughout the training epochs. Finally, the input signal is split into segments of $2400$ samples (equivalent to $50$ ms) to be processed before updating the weights. Considering the network architecture described in Section 3.1, in which a buffer of the $64$ most recent samples is used as input, training layers with states necessitates sequential input-output examples. Consequently, our implementation adopts an input shape of ($B$, $L$, $F$), which is set to ($B$, $1$, $64$), where $B$ denotes the arbitrary batch size, $L$ the sequence length, and $F$ the input dimensionality, often identified as the number of input features. This configuration considers the $64$ input samples as the vector of input features at the current sampling instant.
We adopted the same learning schedule and minibatch size for all models across all datasets. As explained in Section 3.3, we collected the datasets using the same approach and the same input file to minimize the difference among them. However, it is important to highlight that each dataset and modeling task may benefit from specific adjustments to achieve more accurate results. Different datasets and tasks may require slightly different learning rate schedules; an excessively slow or fast learning rate can lead to suboptimal solutions. The number of samples to process before updating the weights referred to as the minibatch size, impacts the frequency of updating weights: larger sizes may result in faster convergence but poorer generalization, while smaller sizes can aid generalization but may require additional epochs.
@@ 4 Results
The evaluation is based on the metrics described in Section 3.2 and presented in Table 4. The MSE loss and the epoch-related epoch resulting in lower validation loss during training are reported in Table 3. The losses and the metrics consider all conditioning scenarios and provide an average error across all combinations of parameters. To further verify the model’s capacity to learn the conditioned effect, we provide a series of plots that compare the target signal with the predictions generated by the four distinct architectures. For these plots, we chose the modeling scenario that we believe to be the most challenging based on the effect parameters used as conditioning factors for the networks. Except for the low-pass filter, this typically involves parameters at the maximum of the selected range since the output signal, under these conditions, is likely to deviate most significantly from the input. Regarding the low-pass filter, the parameters are configured for minimum cutoff and maximum resonance. The plots show the RMSE in Figure  4, and the spectrograms in Figure  5. For datasets including conditioning parameters, the plots showing predictions versus the target are generated for a $2.5$ seconds input signal containing bass, guitar, and drum loop sound. These input signals are part of the respective test sets. To ensure a clearer comparative visualization of the spectral flux, we use only the first $1.25$ seconds.
@@
OD300 LSTM $\boldsymbol{2.46\cdot 10^{-4}}$ $144$ Neutron LSTM $\boldsymbol{2.27\cdot 10^{-3}}$ $4$
Overdrive ED $3.57\cdot 10^{-4}$ $176$ Overdrive ED $2.79\cdot 10^{-3}$ $3$
@@
S6 $3.39\cdot 10^{-4}$ $172$ S6 $2.41\cdot 10^{-3}$ $28$
@@
LRU $\boldsymbol{4.14\cdot 10^{-4}}$ $176$ Equalizer LRU $1.03\cdot 10^{-3}$ $150$
@@
S6 $4.53\cdot 10^{-4}$ $188$ S6 $9.68\cdot 10^{-4}$ $8$
LA-2A LSTM $1.11\cdot 10^{-4}$ $59$ CL 1B LSTM $1.78\cdot 10^{-4}$ $74$
Optical ED $9.54\cdot 10^{-5}$ $90$ Optical ED $\boldsymbol{8.96\cdot 10^{-5}}$ $38$
Compressor LRU $1.14\cdot 10^{-4}$ $146$ Compressor LRU $6.76\cdot 10^{-4}$ $168$
S4D $8.43\cdot 10^{-5}$ $62$ S4D $1.22\cdot 10^{-4}$ $14$
S6 $\boldsymbol{8.25\cdot 10^{-5}}$ $42$ S6 $1.01\cdot 10^{-4}$ $19$
Neutron LSTM $4.07\cdot 10^{-3}$ $46$